OpenStudy (anonymous):

Combine the like terms to simplify the expression: 10xy2 + 5x – 4y3 + 6 – x + 5x2y + 6xy2 – 2 A. 10xy2 + 5x2y – 4xy3 + 4 B. 21xy2 + 4x – 4y3 + 4 C. 16xy2 + 5x2y + 4x – 4y3 + 4 D. 16xy2 + 5x2y + 4 Question 2 Which is 3125 written as a power of a number? A. 55 B. 53 C. 35 D. 65 Question 3 What is the value of 21 + 2a4 – 4b2 when a = -2 and b = 3?

OpenStudy (anonymous):

10xy^2 + 5x - 4y^3 + 6 - x + 5x^2y - 2 Like terms will have exactly the same letters and exponents I willl separate the like terms so you can see.. 10xy^2 -- there is no other like term for this 5x - x = 4x -4y^3 -- there is no other like term for this 6 - 2 = 4 5x^2y -- there is no other like terms for this so we are left with : 10xy^2 + 5x^2y - 4y^3 + 4 ======================= not sure about this one....55^2 = 3025...but not 3125 ======================= Let me refresh the page to do the last one...I am just seeing black diamonds

OpenStudy (anonymous):

21 + 2a^4 - 4b^2 when a = -2 and b = -3 this is just a matter of subbing and solving 21 + 2(-2^4) - 4(-3^2) 21 + 2(-16) - 4(9) 21 - 32 - 36 21 - 68 - 47